How many kgs. of tomato costing $1.20 per kg must be mixed with 12 kg of potato costing $ 0.90 per kg so that 50% gain may be obtained by selling the mixture at $1.10 per kg?
Answer by richwmiller(17219)

1.20*x+0.90*12=1.10(12+x)
1.2*x+10.8=13.2+1.1x
1.2x-1.1x=13.2-10.8
0.1x=2.4
x=24 kg of $1.20 is added to 12 kg of $.90 to get 36 kg at $1.10
check
1.2*24+0.9*12=1.1(12+24)
28.8+10.8=1.1(36)
39.6=39.6
ok